OpenStudy (anonymous):

If A is (-4,3) and B is (-6,-5) what is the length of AB, the slope of it and the midpoint?

OpenStudy (jim766):

length \[\sqrt{(x-x)^2+(y-y)^2}\] slope \[\frac{ y2-y1 }{ x2-x1 }=m\] mid point add the x's and divide by 2, add the y's and divide by 2

OpenStudy (anonymous):

Do you agree that the length is 4 root 13, the slope is 8/9 and the midpoint is (1,-1)?

OpenStudy (anonymous):

or is the slope negative 2/5?

OpenStudy (jim766):

\[\sqrt{(-4- - 6)^2+(3- -5)^2}\] \[\sqrt{(2)^2+(8)^2}\]

OpenStudy (jim766):

\[\sqrt{68}\]

OpenStudy (anonymous):

so 2 root 17 for the length?

OpenStudy (jim766):

slope -5 -3 -8 ----- = --- = 4 -6 - - 4 -2

OpenStudy (jim766):

right 2 root 17

OpenStudy (anonymous):

what about the midpoint?

OpenStudy (jim766):

mid point -4 3 -6 -5 ------ -10 -2 -- -- 2 2 -5, -1 mp

OpenStudy (anonymous):

Thank you very much Jim, I appreciate all the help

OpenStudy (jim766):

yw